# The Internal Model Principle of Time-Varying Optimization

### v2

@@ 4.2 Design of dynamic gradient feedback algorithms
The algorithm uses a Luenberger observer to estimate the exosystem state $\theta(t)$ (cf. line $4$), and a parameter feedback algorithm is then applied to the estimated exosystem state to regulate the gradient to zero (precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 – cf. line $3$ of Algorithm 1). Notice that the algorithm uses the particular choice for $\sigma(\theta)$ being the identity operator (see Remark 12). Note that line $2$ of the algorithm guarantees that the equilibrium of (22) is locally exponentially stable, as required by Theorem 4, and that the existence of such $L$ is guaranteed by Assumption 3.
@@ 1 $n_{c}\leftarrow p$;
2 $L\leftarrow$ any matrix such that $S-LQ$ is Hurwitz;
@@ 4 $F_{c}(z,y)\leftarrow s(z)+L(y-\nabla_{x}f(H_{c}(z),z))$;
Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in{}^{p\times p}$ as in Assumption 2. Further, suppose that $Q$ is such that the unobservable subspace $\mathcal{N}=\cap_{i\geq 0}\ker(QS^{i})$ is the empty set $\mathcal{N}=\emptyset,$ so that Assumption 3 is satisfied; notice that this assumption holds, for example, when $n=p$ and $Q$ is invertible. According to Theorem 5, an optimization algorithm given by
@@
where $A\in{}^{n_{c}\times n_{c}},$ $B_{c}\in{}^{n_{c}\times n},$ $G_{c}\in{}^{n\times n_{c}},$ achieves asymptotic tracking if and only if there exists a linear transformation $\Sigma\in{}^{n_{c}\times p}$ such that:
@@
00 0 $=RG_{c}\Sigma+Q.$ (30b)
Next, we apply Algorithm 1 to design an algorithm that solves this problem. The mapping $H_{c}(\theta)=-R^{\dagger}Q\theta$ zeros the gradient (see Example 2). Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that $S-LQ$ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
@@
which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that $S-LQ$ is Hurwitz. With this choice, conditions (30) are satisfied with $\Sigma=I$ (see Remark 12). Finally, the autonomous model (22) becomes $\dot{z}(t)=(S-LQ)z(t)$, which is asymptotically stable, as required by Theorem 4. $\square$
@@ 4.3 The optimization algorithm as a copy of the exosystem
As shown by (27), the state of optimization algorithms that achieve asymptotic tracking and that of the exosystem are related by a change of coordinates. In this section, we show that $\sigma(\theta)$ is an injective map and thus is a well-posed change of coordinates. For Algorithm 1, we have the following.
@@
It follows from Proposition 7 that, for any optimization method returned by Algorithm 1, there exists an injective mapping that, locally, maps the state of the exosystem into the state of the optimization algorithm.
This property is not exclusive to Algorithm 1; in fact, any algorithm that achieves exact asymptotic tracking possesses the same characteristic, as demonstrated next.
@@
Let the assumptions of Theorem 5 hold and assume that $s(\theta)$ is $C^{\infty}.$ For any $F_{c}(z,y)$ and $G_{c}(z)$ that achieve asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective in $\Theta_{\circ}$$\square$
@@
is not invertible at $\theta=\theta^{\prime}.$ By [53, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
@@ 5 Fidelity of the internal model vs accuracy
In Section 4, we have shown that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem. Departing from this finding, in this section we tackle the following question: how will the algorithm perform when the temporal variability of the problem is known only approximately? For illustration purposes, we focus on the quadratic problem (19) (see Examples 2 and 3). Suppose that an imprecise internal model is available; that is, there exists $\Delta\in{}^{n_{c}\times p}$ such that (30) is modified to777 To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
@@ 6.1 Constrained problems
[…]
@@
For any such saddle-point, if strong duality holds, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero. Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point. It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (10b) is also a saddle-point and strong duality holds, then it is also a solution to the equality-constrained problem. When strong duality does not hold, however, such a saddle-point may not correspond to an optimizer [54, Ch. 5].
@@ 6.2 Discrete-time algorithms
In practice, optimization algorithms are often implemented on digital machines and thus, in these cases, implementing (10) requires discretizations. In these cases, a possible solution is to pursue a design by emulation [55], which consists in converting the continuous-time algorithm (10) into a discrete-time one by approximation (that is, by approximating the time-derivative by a discrete difference using, e.g., Backward/Forward Difference, Tustin, ZOH methods, etc.). Then, the continuous-time guarantees given here can be translated into discrete-time ones (using standard tools from emulation [55] or sampled-data systems [56]), provided that the sampling time is sufficiently small. In these cases, exact tracking established here will continue to hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [55].
An alternative method that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4). For this alternative formulation, the critical trajectories are discrete signals and, in this case, it would be natural to seek a fully-discrete alternative to (10) and pursue an analysis in discrete time. Such a design is beyond the scope of this work and will be left as the focus of future works.
@@ 7.2 Nonlinear example
with parameters $\kappa=7.5$ and $\mu=1.5$, where the time-varying parameter $\theta(t)$ satisfies the exosystem dynamics $\dot{\theta}=S\theta$ where $S=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$. We compare the error $\|x(t)-x^{\star}(t)\|$ over time for several algorithms in Fig. 4. The objective function satisfies the conditions in Theorem 3; for our simulations, we computed numerically a mapping that zeros the gradient locally. In Fig. 4(a), we show the error for our algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (shown in dark to light blue) along with the corresponding rate $e^{-\eta t}$ (dashed black). For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [34, Equations (9) and (20)] using the parameter values in [34, Section VI.A].
@@ 7.3 Application to solve the dynamic traffic assignment problem in transportation
Figure 6: Illustration of the performance of the optimization method synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (7.3). See Section 7.3 for a discussion. The bottom figure shows that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium. Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of $S-LQ.$
@@
Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm – only the frequencies $\omega_{1},\omega_{2}$ are required to be known. To seek a solution to the constrained problem (7.3), consider the Lagrangian function:
[…]
where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers. We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section 6); the inequality constraints in (7.3) have been accounted for by projecting $x(t)$ onto the feasible set. Here, matrix $L$ has been chosen so that the eigenvalues of $S-LQ$ are uniformly distributed in the open real interval $(-1,-2)$. Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [58], and thus the problem admits a unique critical point that is a saddle point. It follows that our algorithm is guaranteed to converge to a minimizer of (7.3). Simulations for this problem are presented in Fig. 6. The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig. 6-second figure) and $y(t)\rightarrow 0$ (up to machine precision) which implies that the algorithm successfully computes a critical point (see Fig. 6-third figure). Finally, the bottom figure of Fig. 6 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
@@ 8 Conclusions
We showed that the problem of designing algorithms for time-varying optimization problems can be reformulated as a nonlinear multivariate output regulation problem. This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to functional evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem. On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zero the gradient. Importantly, our results show that asymptotic tracking can be achieved under more relaxed assumptions than what is normally imposed in the literature. Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem. This work opens the opportunity for several directions of future work, including the relaxation of the convexity and smoothness assumptions, an investigation of discrete-time problems, and applications to feedback optimization.
@@
A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37) satisfying $x(0)=x_{\circ}$ satisfiest $x(t)\in S$ for all $t\in(t_{1},t_{2})$. $\square$
@@
a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only. The remainder of this section is devoted to formalizing this fact.
@@
Assume that $n^{+}=0$. There exists a neighborhood $V\subset{}^{n^{0}}$ of 00 and a class $C^{k-1}$ mapping $\pi:V\rightarrow{}^{n^{-}}$ such that the set
$S=\{(y,\theta)\in{}^{n^{-}}\times V:y=\pi(\theta)\},$
@@
$\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$square

### v3

@@ 4.2 Design of dynamic gradient feedback algorithms
The algorithm uses a Luenberger observer to estimate the exosystem state $\theta(t)$ (cf. line $4$), and a parameter feedback algorithm is applied to the estimated exosystem state to regulate the gradient to zero; precisely, $G_{c}(z)$ is designed following the approach of Theorem 1 (see line $3$ of Algorithm 1). Notice that the algorithm uses the particular choice for $\sigma(\theta)$ being the identity operator (see Remark 12). Note that line $2$ of the algorithm guarantees that the equilibrium of (22) is locally exponentially stable, as required by Theorem 4, and that the existence of such $L$ is guaranteed by Assumption 3.
@@ 1 $n_{c}\leftarrow p$;
2 $L\leftarrow$ any matrix such that SLQS-LQ is Hurwitz;
@@ 4 $F_{c}(z,y)\leftarrow s(z)+L(y-\nabla_{x}f(H_{c}(z),z))$;
Consider the quadratic problem studied in Example 2, and assume that the exosystem follows the linear model $\dot{\theta}=S\theta$ for some matrix $S\in\real^{p\times p}$ as in Assumption 2. Further, suppose that $Q$ is such that the unobservable subspace $\mathcal{N}=\cap_{i\geq 0}\ker(QS^{i})$ is the empty set $\mathcal{N}=\emptyset,$ so that Assumption 3 is satisfied; notice that this assumption holds, for example, when $n=p$ and $Q$ is invertible. According to Theorem 5, an optimization algorithm given by
@@
where $A\in\real^{n_{c}\times n_{c}},$ $B_{c}\in\real^{n_{c}\times n},$ $G_{c}\in\real^{n\times n_{c}},$ achieves asymptotic tracking if and only if there exists a linear transformation $\Sigma\in\real^{n_{c}\times p}$ such that:
@@
$0$ $=RG_{c}\Sigma+Q.$ (30b)
Next, we apply Algorithm 1 to design an algorithm that solves this problem. The mapping $H_{c}(\theta)=-R^{\dagger}Q\theta$ zeros the gradient (see Example 2). Following Algorithm 1, we set $n_{c}=p$, choose matrix $L$ such that SLQS-LQ is Hurwitz, define $G_{c}(z)=H_{c}(z)=-R^{\dagger}Qz$, and then select $A_{c}$ and $B_{c}$ such that $F_{c}(z,y)=A_{c}z+B_{c}y$ satisfies:
@@
which gives $A_{c}=S$ and $B_{c}=L.$ In summary, a dynamic gradient-feedback algorithm is (29) with $A_{c}=S$, $B_{c}=L$, and $G_{c}=-R^{\dagger}Q$, where $L$ is any matrix such that SLQS-LQ is Hurwitz. With this choice, conditions (30) are satisfied with $\Sigma=I$ (see Remark 12). Finally, the autonomous model (22) becomes $\dot{z}(t)=(S-LQ)z(t)$, which is asymptotically stable, as required by Theorem 4. $\square$
@@ 4.3 The optimization algorithm as a copy of the exosystem
As shown in the previous section, the state of the optimization algorithm and that of the exosystem must be related, at all points of the limit set, by (27). In this section, we show that $\sigma(\theta)$ is injective, and therefore defines a well-posed change of coordinates. We begin by establishing the following result.
@@
It follows from Proposition 7 that, for any optimization method returned by Algorithm 1, there exists an injective mapping that, locally, maps the state of the exosystem into the state of the optimization algorithm. This property is not exclusive to Algorithm 1; in fact, any algorithm that achieves exact asymptotic tracking possesses the same characteristic, as demonstrated next.
@@
Let the assumptions of Theorem 5 hold, and assume that $s(\theta)$ is of class $C^{\infty}$. Then, for any algorithm of the form (10) that achieves asymptotic tracking, there exists a neighborhood $\Theta_{\circ}\subset\Theta$ of the origin such that $\sigma(\theta)$ is injective on $\Theta_{\circ}\square$
@@
is not invertible at $\theta=\theta^{\prime}.$ By [54, Thm. 3.13], the system (2) is not weakly observable, thus violating Assumption 3. ∎
@@ 5 Fidelity of the internal model vs accuracy
In Section 4, we showed that a necessary condition for exact asymptotic tracking is that the algorithm incorporates an internal model of the temporal variability of the problem. Departing from this finding, in this section we tackle the following question: how will the algorithm perform when the temporal variability of the problem is known only approximately? For illustration purposes, we focus on the quadratic problem (19) (see Examples 2 and 3). Suppose that an imprecise internal model is available; that is, there exists $\Delta\in\real^{n_{c}\times p}$ such that (30) is modified to [^7] To explicitly model uncertainty in the matrix $S,$ equation (32) could be rewritten as $\Sigma(S+\Delta_{s})=A_{c}\Sigma.$ By comparison, the two approaches are equivalent with the substitution $\Delta=\Sigma\Delta_{s}.$ :
@@ 6.1 Constrained problems
where $\lambda_{i}$ is the Lagrange multiplier associated with the $i\textsuperscript{th}$ equality constraint. A pair $(x,\lambda)\in\real^{n}\times\real^{m}$ is said to be a saddle-point of the Lagrangian if, for all $(\bar{x},\bar{\lambda})\in\real^{n}\times\real^{m}$,
@@
For any such saddle-point, if strong duality holds, $x$ is primal optimal, $\lambda$ is dual optimal, and the optimal duality gap is zero. Moreover, the gradient of the Lagrangian (assuming it exists) is zero at any saddle-point. It follows from the derivations in the previous sections that the gradient-feedback and parameter-feedback algorithms can be directly applied to seek a stationary point of the Lagrangian function by replacing (in (1)) the variable $x$ with the extended decision variable $\tilde{x}=(x,\lambda)$ and by letting $f(\tilde{x},\theta)=L(x,\lambda,\theta).$ Notice that, if the critical point computed by (10b) is also a saddle-point and strong duality holds, then it is also a solution to the equality-constrained problem. When strong duality does not hold, a saddle-point may not correspond to an optimizer [55, Ch. 5].
@@ 6.2 Discrete-time algorithms
In practice, optimization algorithms are often implemented on digital machines and thus require discretization of (10). It is worth noting that this scenario corresponds to designing a discrete-time algorithm to track a continuous-time optimizer (cf. (1) and (5)). In these cases, a possible solution is to pursue a design by emulation [56], which consists in converting the continuous-time algorithm (10) into a discrete-time one by approximation (that is, by approximating the time-derivative by a discrete difference using, e.g., Backward/Forward Difference, Tustin, ZOH methods, etc.). Then, the continuous-time guarantees given here can be translated into discrete-time ones using standard tools from emulation [56], provided that the sampling time is sufficiently small. In these cases, exact tracking established above will hold at the discrete time instants, while tracking will be inexact in-between sampling times and the error can be bounded in terms of the sampling frequency using the explicit bounds in [56].
An alternative approach that could be pursued is that of discretizing the temporal variability $\theta(t)$, thus leading to a discrete-time optimization problem as in (4). For this alternative formulation, the critical trajectories are instead discrete signals and, in this case, it would be natural to seek a discrete algorithm, in place of (10), and pursue an analysis in discrete time. Such a design is beyond the scope of this work and is left as the focus of future works.
@@ 7.2 Nonlinear example
with parameters $\kappa=7.5$ and $\mu=1.5$, where the time-varying parameter $\theta(t)$ satisfies the exosystem dynamics $\dot{\theta}=S\theta$ where $S=\left[\begin{smallmatrix}0&1\\ 1&0\end{smallmatrix}\right]$. We compare the error $\|x(t)-x^{\star}(t)\|$ over time for several algorithms in Fig. 4. The objective function satisfies the conditions in Theorem 3; for our simulations, we computed numerically a mapping that zeros the gradient locally. In Fig. 4(a), we show the error for our algorithm with observer pole locations $(-\eta,-2\eta)$ for $\eta\in\{1,5,20\}$ (shown in dark to light blue) along with the corresponding rate $e^{-\eta t}$ (dashed black). For comparison, we also show the error of the prediction-correction (PC) algorithm (8) along with two of its finite-time (FT) variants [35, Equations (9) and (20)] using the parameter values in [35, Section VI.A].
@@ 7.3 Application to solve the dynamic traffic assignment problem in transportation
Figure 6: Illustration of the performance of the optimization method synthesized using Algorithm 1 to solve the dynamic traffic assignment problem (7.3). See Section 7.3 for a discussion. The bottom figure shows that, from a zero initial condition, the algorithm is capable of computing a Wardrop equilibrium in about $1.5$ hours, and then of tracking this equilibrium. Notice that, a faster rate of convergence could be obtained by shifting the eigenvalues of SLQ.S-LQ.
@@
Notice that knowledge of $\theta_{0},\theta_{1},\theta_{2},\phi_{1},\phi_{2}$ is not required to synthesize the optimization algorithm—only the frequencies $\omega_{1},\omega_{2}$ are required to be known. To seek a solution to the constrained problem (7.3), consider the Lagrangian function:
$L(x,\lambda,\theta(t)):=\sum_{i\in\mathcal{E}}\int_{0}^{x_{i}}\ell_{i}(s)ds\\ +\sum_{v\in\mathcal{V}}\lambda_{v}\biggl(\delta_{v}(\theta(t))-\sum_{j\in\mathcal{E}:j^{-}=v}x_{j}+\sum_{j\in\mathcal{E}:j^{+}=v}x_{j}\biggr),$
where $\lambda:=(\lambda_{1},\dots,\lambda_{|\mathcal{V}|})$ is the vector of Lagrange multipliers. We applied Algorithm 1 to the optimization problem (1) with $f(\tilde{x},\theta)=L(x,\lambda,\theta)$ (see Section 6); the inequality constraints in (7.3) have been accounted for by projecting $x(t)$ onto the feasible set. Here, matrix $L$ has been chosen so that the eigenvalues of SLQS-LQ are uniformly distributed in the open real interval (1,2)(-1,-2). Notice that, since the latencies are strictly increasing, the Lagrangian is strongly convex-strongly concave [58], and thus the problem admits a unique critical point that is a saddle point. It follows that our algorithm is guaranteed to converge to a minimizer of (7.3). Simulations for this problem are presented in Fig. 6. The simulation shows that $z(t)\rightarrow\theta(t)$ as $t\rightarrow\infty$ (see Fig. 6-second figure) and $y(t)\rightarrow 0$ (up to machine precision) which implies that the algorithm successfully computes a critical point (see Fig. 6-third figure). Finally, the bottom figure of Fig. 6 illustrates the rate of convergence of the algorithm, which, as expected, is governed by the placement of the eigenvalues of the observer.
@@ 8 Conclusions
We showed that the problem of designing algorithms for time-varying optimization problems can be reformulated as a nonlinear multivariate output regulation problem. This connection allowed us to prove the internal model principle of time-varying optimization, which states that (when the algorithm has access only to function evaluations of the gradient) asymptotic tracking can be achieved only if the algorithm incorporates a reduplicated model of the temporal variability of the problem. On the other hand, when the time-varying parameters are measurable, asymptotic tracking can be achieved by seeking a mapping that zeros the gradient. Importantly, our results show that asymptotic tracking can be achieved under more relaxed assumptions than what is normally imposed in the literature. Moreover, our algorithm structure is novel in the literature, and it relies on the use of an observer for the temporal variability of the problem. This work opens the opportunity for several directions of future work, including the relaxation of the convexity and smoothness assumptions, an investigation of discrete-time problems, and applications to feedback optimization.
@@
A $C^{k}$ manifold $S$ of $U$ is locally invariant for (37) if, for each $x_{\circ}\in S$, there exists $t_{1}<0<t_{2}$ such that the integral curve $x(t)$ of (37), such that $x(0)=x_{\circ}$, satisfies $x(t)\in S$ for all $t\in(t_{1},t_{2})$. $\square$
@@
a curve $y=\pi(\theta)$ is an invariant manifold for (38) if the solution of (38) with $\theta(0)=\theta_{\circ}$ and $y(0)=\pi(\theta_{\circ})$ lies on the curve $y=\pi(\theta)$ for $t$ in a neighborhood of $0.$ The notion of invariant manifold is useful as, under certain assumptions, it allows us to reduce the analysis of (37) to the study of a reduced system in the variable $\theta$ only. The remainder of this section is devoted to formalizing this fact.
@@
Assume that $n^{+}=0$. There exists a neighborhood $V\subset\real^{n^{0}}$ of 0 and a class $C^{k-1}$ mapping $\pi:V\rightarrow\real^{n^{-}}$ such that the set
$S=\{(y,\theta)\in\real^{n^{-}}\times V:y=\pi(\theta)\},$
@@
$\|y(t)-\pi(\theta(t))\|\leq Me^{-Kt}\|y(0)-\pi(\theta(0))\|.$ $\square$